### v1

@@ 6.1 Low spectral indices or low brightnesses?
To alleviate the effect of different samplings of interferometer observations, we re-clean the 1.3 and 3 mm continuum images to achieve a common beam, and re-run the getsf to obtain common-beam versions of dbcat. The common-beam dbcat is specifically used for inter-band spectral index calculation (see more details in Appendix E). We note that in our published dbcat (Table 6), the source extractions are still from the images with original beams, because common-beam dbcat will miss detections of blended sources due to its coarser resolution.
@@
The spectral index is calculated from common-beam dbcat using Eq. (1), which is plotted against the beam-filling peak intensity at 1.3 mm, , which represents the peak intensity when the beam filling factor is unit. For sources with deconvolved sizes smaller than the beam size, a beam filling factor correction of is applied to the measured peak intensities to obtain . It is convenient to compare with each other as well as to the modified blackbody emission in the later discussion. All the dbcat sources with reliable spectral indices are shown by gray data points with error bars, and the KDE of their distribution is outlined by gray contours.
@@
We highlight the KDE distribution of in the top panel of Fig. 6, where it shows a mean value of and a median value of for all the three clouds. As a comparison, we calculate the track on the - plane of modified blackbody (MBB) emission in the form of Eq. (2) and with a given temperature (see details in Appendix F), which is further referred as isothermal track, or Eq. (20). If data points follow the isothermal track with temperature from 10 to 30 K and , then their could be calculated as color-coded KDEs in the top panel of Fig. 6. The observed distribution, however, systematically deviates from the MBB emission with the commonly-assumed and with the assumed beam filling factor of unity. Such deviations account for more than 70% of the detections.
@@
We now interpret the deviation in two ways, which guides us to propose different scenarios and hypotheses in the following sections. One is that the dust emissivity spectral index is as low as , so that the optically thin dust emission could show , which is referred to as the “low spectral indices” case. The other is that the peak intensity or brightness temperature could be lower than what MBB expects. As the same way that we did for above, can be also predicted for given spectral indices and dust temperatures. We show them by the color-coded KDEs in the right panel of Fig. 6. These distributions, which peak at mJy, cannot account for the significantly lower observed values of , which peak around 1 mJy beam-1. We refer to this as the “low brightness temperature” case.
@@ 6.2 Possible explanations
One possibility to explain the observed low spectral indices and low brightnesses is that the flux densities are dominated by spatially compact and optically thick objects that are severely beam diluted. For example, the flux densities may be dominated by embedded Class 0/I Young Stellar Objects (YSOs). In Sect. 6.2.1, we explore what properties of such compact sources may be in order to explain the observed spectral indices and brightnesses.
Meanwhile, the presence of 0.1 mm sized dust grains can also lead to attenuated and lower spectral indices, due to the high dust scattering opacity associated with larger grains. In case that flux densities are not necessarily dominated by embedded Class 0/I YSOs in some sources, in Sect. 6.2.2, we explore the extent of grain growth needed to explain the observed spectral indices and brightnesses. Finally, sources associated with free-free emission typically exhibit relatively lower spectral indices, as discussed in Sect. 6.2.3.
We note that the three possibilities are not mutually exclusive. For example, with the presence of large dust grains, embedded Class 0/I YSOs with larger angular scales (i.e., not severely beam diluted) can be consistent with the observations, as the attenuation due to both dust scattering and beam dilution helps explain the observed low brightnesses.
@@ 6.2.1 Beam diluted Class 0/I YSOs
The hypothesis is motivated by recent ALMA long-baseline observations towards several CMZ clouds, revealing a large population of hundred-au scale compact sources (Budaiev et al. 2024; Zhang et al. 2025a). Combining with the “low brightness” case introduced in Sect. 6.1 and the small sizes discussed in Sect. 5.1, our hypothesis is that these low- sources could be unresolved Class 0/I YSOs, whose brightness temperature is significantly beam-diluted when being observed with a large beam size.
@@
To verify the possibility, we retrieve the 1.3 mm continuum sources from the ALMA Survey of Orion Planck Galactic Cold Clumps (ALMASOP; Dutta et al. 2020) with 140 au resolution. A total of 66 sources have been classified as Class 0/I YSOs and four are prestellar core candidates. As indicated in Fig. 7 (a), the sizes of Class 0/I YSOs are between 50 and 200 au (16% and 85% percentile for ), while the prestellar cores are usually as large as au, which is consistent with other well-resolved prestellar cores (e.g., Caselli et al. 2022). We adopt the dust models of 10-micron-sized grains (details in Sect. 6.2.2) and obtain their isothermal tracks with optically thick diameter of 50, 100, and 200 au. The isothermal tracks are spanned as shaded regions with a temperature range of 30–100 K, shown with gray, pink, and blue colors in Fig. 8. These models can cover the observed data points very well. In this case, a large fraction of the observed flux densities can be contributed by the embedded, optically thick low-mass Class 0/I YSOs that are much smaller than the sizes of our synthesized beams.
@@
To further prove the correspondence between our low-brightness objects and the low-mass Class 0/I YSOs that have been observed in Orion, we perform simulated DUET 1.3-mm observations of the ALMASOP sources by using the simalma task in CASA. As shown in Fig. 7 (b), the integrated fluxes of the ALMASOP Class 0/I objects are primarily between 5 mJy and 120 mJy (16% and 85% percentile). So we choose four representative ALMASOP targets, G210.9719.33S2, G208.6819.20N3, G208.6819.20N1, and G208.6819.20N2, for simulated observations. The first three are identified as Class 0/I YSO candidates, with integrated flux covering from 10 to 100 and to 800 mJy. The last one G208.6819.20N2 is classified as a prestellar core with a 150-au extremely dense nucleus (Hirano et al. 2024).
@@
Data from all three arrays combined (C-5, C-2, and ACA) are used to achieve the highest angular resolution as well as to recover diffuse emission and large-scale structures especially for the prestellar core G208.6819.20N2. Details of the ALMASOP observations, array configuration combination, and imaging can be found in Dutta et al. (2020). We extract the emission above for each image as sky models and put them at the same distance as our studied CMZ clouds, which are shown in Cols. 1 and 3 of Fig. 9. We set simulated observations with the same array configurations (C40-5 and C40-3 in Cycle 4) and the same integration time (270 and 90 seconds) as the actual DUET Band 6 observations. The simulated uv data are cleaned, and the achieved rms noises of cleaned images are 40 Jy beam-1, which is very close to those of our DUET images.
@@
Could these sub-mJy sources be prestellar cores? Our simulated observation of G208.6819.20N2 suggests this possibility. As shown in the bottom panel of Col.4, when placed at the distance of the CMZ, G208.6819.20N2 is detected above , with a peak intensity reaching 0.27 mJy beam-1. Given that G208.6819.20N2 is an extreme case—the brightest and densest prestellar core in the ALMASOP sample—prestellar cores could plausibly account for detections in the range of 0.1 to 0.3 mJy beam-1. Examples might include SgrC-1mm#208 in the SgrC “Bridge” and SgrC-1mm#211 in the SgrC ”Main” region. However, the possibility that even brighter sources ( mJy beam-1) could also be prestellar cores remains to be verified, as the free-fall time in such cases might be even shorter, making these objects exceptionally rare.
There is more evidence of our hypothesis from recent higher resolution observations towards the CMZ clouds. For example, there are 199 bright and compact millimeter sources with a median size of 370 au in three massive star-forming clumps in the Sgr C and the 20 km s-1 cloud, which are suggested to be candidates of protostellar envelopes or disks (Zhang et al. 2025a). With a bit lower spatial resolution of 500–700 au, Budaiev et al. (2024) have also identified more than 200 sources with similar sizes in the Sgr B2 region, which are argued to be Class 0/I YSO candidates with actively accreting protostars and rotation-supported envelopes or disks.
@@ 6.2.2 Large dust grains
Here we discuss the possibility of interpreting our observations as the presence of mm/cm-sized large grains on 103 au scales. Large mm/cm-sized grains generally have lower millimeter absorption opacity spectral indices and therefore lower . Theoretically, Draine (2006) shows that for typical dust in the interstellar medium, at 1.3 mm if the power-law size distribution of the dust grains ( with ) extends to sizes of mm. Recently, there are multi-band observations in protostellar cores in Orion attribute the millimeter shallow spectral indices to larger grain sizes (Schnee et al. 2014; Nozari et al. 2025).
When the dust emission is optically thin, (and thereby possible grain growth in the (sub)millimeter-sized regime) can be inferred from measuring the (sub)millimeter spectral indices (Equation 16). However, dense cores and the embedded Class 0/I objects are not necessarily optically thin at (sub)millimeter wavelengths (e.g., Li et al. 2017; Galván-Madrid et al. 2018). In these cases, it is insufficient to only consider the dust absorption opacity and its spectral index . A self-consistent diagnostic of dust grain growth also needs to take dust scattering opacity into consideration (Liu 2019). Since forward-scattered photons are not effectively scattered and therefore are not relevant in the radiative transfer, it is only necessary to consider the effective scattering opacity, , which excludes the cross-section in the forward scattering. In this case, the total extinction opacity () is expressed by
@@
The relative importance between and can be characterized by dust albedo . In the cases with , spectra of optically thick dust emission is identical to the Planck function . When , the effect of dust scattering starts to become noticeable in the dust spectra when (Birnstiel et al. 2018). For most of the probable dust compositions (except for graphitic carbonaceous dust), this occurs in the Rayleigh-Jeans limit (i.e., ). First, dust scattering leads to attenuation of the observed intensity. Second, it changes the color (i.e., spectral indices), which depends on how the relative importance between and varies with frequency. Spectral indices are lowered (as compared to the dust with zero scattering opacity) in the frequency range where albedo is increasing with frequency, which was referred to as anomalous reddening (Liu 2019); conversely, spectral indices become larger in the frequency range where albedo is decreasing with frequency.
@@
To explore these effects and compare with our observations, we produced a grid of dust spectra based on various assumptions of dust temperature, dust column density, and . The lowest and highest temperatures we considered are 6 K and 48 K, which correspond to the typical ISM temperature lower limit and Rayleigh-Jeans limits, respectively. When evaluating dust spectral profiles, we adopt the formulation for an isothermal dust slab provided in Birnstiel et al. (2018) that is based on the Mie theory and the Henyey–Greenstein scattering approximation. In addition, we adopt the dust opacity tables provided in Birnstiel et al. (2018) which assume that dust grains are spherical and compact and are composed of water ice, refractory organics, troilite, and astrophysical silicates. When evaluating the size-averaged opacities, we assume a power-law grain size distribution with between the minimum grain size ( 0.1 m) and the assumed maximum size . The size-averaged opacity is not sensitive to the assumption of . For each grain size and a given temperature, surface densities start from to 30 g cm-2 to form an isothermal track from optically thin through thick cases.
@@
For micron-sized ( mm) dust grains, absorption dominates over scattering at millimeter bands. As shown by the gray region in Fig. 10, the spectral index is close to the case of Eq. (16), so the highest possible spectral index is 3.8. As discussed in Sect. 6.1, with mm, a large amount of sources with low spectral indices cannot be explained.
For mm-sized ( mm) dust grains, in the optically thin regime, the left and right boundaries of spectral indices shift toward higher values as compared to the case of mm. This is because albedo decreases with frequency at 100–200 GHz when mm (see Figures 2 and 3 of Liu 2019), which leads to higher spectral indices in the regime of . As shown by the pink shaded region, the 1 mm sized grain scenario can cover the “high bump” that includes detections with . The high- bump feature actually serves as an observational hint for the existence of intermediate-sized grains in astrophysical environments. In the optically thick regime, the intensity attenuation is noticeable as compared with the case of mm, which better explains the data points with low and low .
@@
The origin of mm/cm-sized grains is uncertain. A typical envelope density of cm-3 is insufficient to form grains of this size. Only at densities of cm-3 would the formation of such large grains be feasible (Hirashita & Omukai 2009; Navarro-Almaida et al. 2024). According to coagulation models, mm-sized grains observed in protostellar envelopes cannot form locally within the envelopes themselves (Wong et al. 2016; Silsbee et al. 2022). In contrast to the inefficient dust growth in the envelope, dust can quickly grow to the order of mm in disks, with a typical timescale of years, thanks to higher densities in disks even in Class 0/I YSOs. Numerical calculations have shown that large grains decoupled from gas can be ejected from disks through outflows due to the centrifugal force, enriching dust in envelopes (Wong et al. 2016; Tsukamoto et al. 2021). The timescale is much shorter than the lifetime of YSOs, therefore allowing large grains to survive in envelopes.
@@ 6.2.3 Free-free emission
Massive protostars can ionize the surrounding gas, which produces bremsstrahlung (or free-free) emission by electronic braking. The free-free emission spectral energy distribution (SED) shows a weak frequency dependence in the optically thin regime at short wavelengths (Kurtz 2005), which can flatten the millimeter SED () by MMB.
To explore the free-free emission contamination, we use the archived JVLA 1.3 cm observations with representative frequency of 23 GHz (Lu et al. 2019b) with rms noise of 0.05–0.15 mJy beam-1 as shown by yellow contours in Fig. 13. There exists extended filamentary 1.3 cm emission, which is suggested to have non-thermal origins (Clark & Caswell 1976; Ho et al. 1985; Lu et al. 2003; Meng et al. 2019; Bally et al. 2024). For example, the one at the northeast corner of the 20 km s-1 cloud has a spectral index of 0.4 (Ho et al. 1985; Padovani et al. 2019), which cannot be ascribed to free-free emission. In our analyses, two elongated structures (the aforementioned one in the 20 km s-1 cloud and one in cloud e) are excluded, because they have non-thermal origins. A spectral index of is assumed for the free-free emission in order to extrapolate from 1.3 cm to millimeter wavelength in the optically thin regime. We discuss the free-free contamination of all the associated sources in Appendix G.
However, these cases can only explain the low spectral indices of a limited number of sources that are associated with detectable radio emission. Below the 0.1 mJy sensitivity at 1.3 cm, there could still be some undetected UCHii regions. Below we quantitatively estimate how much the centimeter continuum contamination would be in order to significantly reduce the spectral index. Considering an optically thin source with a spectral index of 3.8, let the flux at 1.3 mm from pure dust emission be , the expected flux at 3 mm would then be . Taking the free-free emission flux at both 1.3 and 3 mm into account 555Free-free emission has % flux difference at 1.3 and 3 mm due to the weak frequency dependence ()., the differentiation of the spectral index can be rewritten from Eq. (1) as,
@@
Given that more than 70% of the sources exhibit low spectral indices, it is difficult to attribute them entirely to contamination by free-free emission. First, such a large number of massive protostars within a single cluster would conflict with expectations from the IMF, which predicts a decreasing population at the high-mass end. Second, low-mass protostars are unlikely able to contribute to the needed free-free emission either. In nearby molecular clouds (e.g., Corona Australis, Taurus, Ophiuchus, and Perseus; Liu et al. 2014; Dzib et al. 2015; Coutens et al. 2019; Tychoniec et al. 2018), free-free emission from low-mass Class 0/I objects typically exhibits flux densities of only a few mJy or less at centimeter wavelengths. When scaled to the Galactic center, these emissions would become fainter by approximately three orders of magnitude, therefore contributing less than what could result in significant contamination at millimeter bands. To firmly exclude the possibility, the sensitivity needs to be at least (sub-)Jy. Indeed, a recent deep JVLA survey of the high-mass star-forming region G14.2250.50 ( 2 kpc) indicates that free-free emission from low-mass Class 0/I YSOs in distant star-forming regions is challenging to detect with current facilities (Díaz-Márquez et al. 2024).

### v2

@@ 6.1 Low spectral indices versus low brightnesses
To remove the effect of different samplings between 1.3 mm and 3 mm ALMA observations, we re-cleaned the 1.3 and 3 mm continuum images to achieve a common beam (see more details in Appendix E). We reran the getsf to obtain common-beam versions of dbcat. The common-beam dbcat is specifically used for inter-band spectral index calculations. We note that in our published dbcat (Table 6), the source extractions are still from the images with original beams, because common-beam dbcat will miss detections of blended sources due to its coarser resolution.
@@
The spectral indices were calculated from common-beam dbcat using Eq. (1). We define beam-filling peak intensity, , which represents the peak intensity when the beam filling factor is unit. For sources with deconvolved sizes smaller than the beam size, a beam filling factor of is divided from the measured peak intensities to obtain . The correction is convenient for comparison in later discussion.
@@
In Fig. 6, the dbcat sources with reliable spectral indices are displayed in the - plane. The mean and median values of are and , respectively. As a comparison, we calculated how MBB (i.e., Eq. (2)) behaves on the - plane with a given temperature (see deduction in Appendix F), which is further referred to as the isothermal track in form of Eq. (20). To demonstrate the deviation from MBB emission, is calculated for the given of observed sources following isothermal track with a temperature from 10 to 30 K and . The distribution of these values are shown by color-coded KDEs in the top panel of Fig. 6. Significantly, the observed distribution deviates from the MBB emission with the commonly assumed and with the assumed beam filling factor of unity. Such deviations account for more than 70% of the detections. Since we observed lower than MBB expects, we call it ‘low spectral index” case.
@@
We now interpret the deviation in another way. In the right panel of Fig. 6, the observed spans a wide range from 0.1 to 100 mJy but peaks at 1 mJy. Following the MBB isothermal track with given low spectral indices by observation, the expected is given in color-coded KDEs. These distributions, which peak at mJy, cannot account for the observed values which are significantly lower. We refer to this as the “low brightness temperature” case. The two cases discussed above guides us to propose different scenarios and hypotheses in the following sections.
@@ 6.2 Possible explanations
One hypothesis to explain the observed low spectral indices and low brightnesses is that the flux densities are dominated by spatially compact and optically thick objects that are severely beam diluted. For example, the flux densities may be dominated by embedded class 0/I YSOs. In Sect. 6.2.1, we explore what properties of such compact sources may be in order to explain the observed spectral indices and brightnesses. On the other hand, the presence of mm sized dust grains can also lead to attenuated and lower spectral indices, due to the high dust scattering opacity associated with larger grains. In case flux densities are not necessarily dominated by embedded class 0/I YSOs in some sources, in Sect. 6.2.2 we explore the extent of grain growth needed to explain the observed spectral indices and brightnesses. Finally, sources associated with free-free emission typically exhibit lower spectral indices, as is discussed in Sect. 6.2.3.
We note that the three hypotheses are not mutually exclusive. For example, with the presence of large dust grains, embedded class 0/I YSOs with larger angular scales (i.e., not severely beam diluted) can be consistent with the observations, as the attenuation due to both dust scattering and beam dilution helps explain the observed low brightnesses.
@@ 6.2.1 Beam diluted class 0/I YSOs
The hypothesis is motivated by recent ALMA long-baseline observations of CMZ clouds, revealing a large population of hundred-au scale compact sources (Budaiev et al. 2024; Zhang et al. 2025a). Combining with the low brightness (Sect. 6.1) and the small sizes discussed in Sect. 5.1, our hypothesis is that these low- sources could be unresolved class 0/I YSOs, of which the brightness temperature is significantly beam-diluted when being observed with a large beam size.
@@
To verify the hypothesis, we retrieve the 1.3 mm continuum sources from the ALMA Survey of Orion Planck Galactic Cold Clumps (ALMASOP; Dutta et al. 2020) with 140 au resolution. A total of 66 sources have been classified as class 0/I YSOs and four are prestellar core candidates. As is indicated in Fig. 7 (a), the sizes of class 0/I YSOs are between 50 and 200 au (16% and 85% percentile for ), while the prestellar cores are usually as large as au, which is consistent with other well-resolved prestellar cores (e.g., Caselli et al. 2022). We adopted the dust models of 10-micron-sized grains (see description in Sect. 6.2.2) and obtained their isothermal tracks with an optically thick diameter of 50, 100, and 200 au. The isothermal tracks with temperature range of 30–100 K are shown in shaded regions, shown with gray, pink, and blue colors in Fig. 8. These models can cover the observed data points very well. In this case, a large fraction of the observed flux densities can be contributed by the embedded, optically thick low-mass class 0/I YSOs that are much smaller than the sizes of our synthesized beams.
@@
We further perform simulated DUET 1.3-mm observations of the ALMASOP sources by using the simalma task in CASA. As is shown in Fig. 7 (b), the integrated fluxes of the ALMASOP class 0/I objects are primarily between 5 mJy and 120 mJy (16% and 85% percentile). So we choose four representative ALMASOP targets, G210.9719.33S2, G208.6819.20N3, G208.6819.20N1, and G208.6819.20N2, for simulated observations. The first three are identified as class 0/I YSO candidates, with integrated flux covering from 10 to 100 and to 800 mJy. The last one G208.6819.20N2 is classified as a prestellar core with a 150-au extremely dense nucleus (Hirano et al. 2024).
@@
Data from all three arrays combined (C-5, C-2, and ACA) are used to achieve the highest angular resolution as well as to recover diffuse emission and large-scale structures especially for the prestellar core G208.6819.20N2. Details of the ALMASOP observations, array configuration combination, and imaging can be found in Dutta et al. (2020). We extract the emission above for each image as sky models and put them at the same distance as our studied CMZ clouds, which are shown in Cols. 1 and 3 of Fig. 9. We set simulated observations with the same array configurations (C40-5 and C40-3 in Cycle 4) and the same integration time (270 and 90 seconds) as the actual DUET band 6 observations. The simulated visibility data are cleaned, and the achieved rms noises of cleaned images are 40 Jy beam-1, consistent with our DUET images.
@@
Our simulated observation of G208.6819.20N2 also proves the possibility of those sub-milli-Jansky sources to be prestellar cores. In the Col. 4 bottom panel, G208.6819.20N2 is detected above , with a peak intensity reaching 0.27 mJy beam-1, when placed at the distance of the CMZ. Given that G208.6819.20N2 is an extreme case—the brightest and densest prestellar core in the ALMASOP sample—prestellar cores could plausibly account for detections in the range of 0.1 to 0.3 mJy beam-1. Examples might include SgrC-1mm#208 in the SgrC “Bridge” and SgrC-1mm#211 in the SgrC “Main” region. However, the possibility that even brighter sources ( mJy beam-1) could also be prestellar cores remains to be verified, as the free-fall time in such cases might be even shorter, making these objects exceptionally rare.
There are more supports from recent higher resolution observations towards the CMZ clouds. For example, there are 199 bright and compact millimeter sources with a median size of 370 au in three massive star-forming clumps in the Sgr C and the 20 km s-1 cloud, which are suggested to be candidates of protostellar envelopes or disks (Zhang et al. 2025a). With a bit lower spatial resolution of 500–700 au, Budaiev et al. (2024) have also identified more than 200 sources with similar sizes in the Sgr B2 region, which are argued to be class 0/I YSO candidates with actively accreting protostars and rotation-supported envelopes or disks.
@@ 6.2.2 Large dust grains
We discuss the hypothesis of millimeter- and/or centimeter-sized (mm/cm-sized) large dust grains at the scale of thousands of au. Large grains generally have lower millimeter absorption opacity spectral indices and therefore lower . Theoretically, Draine (2006) shows that for typical dust in the interstellar medium, at 1.3 mm if the power-law size distribution of the dust grains ( with ) extends to sizes of mm. Recently, there are multiband observations in protostellar cores in Orion attribute the millimeter shallow spectral indices to larger grain sizes (Schnee et al. 2014; Nozari et al. 2025).
We apply a self-consistent dust grain model which considers both absorption and scattering (Liu 2019). In the model, the effective scattering opacity is , which excludes the cross-section in the forward scattering. So, the total opacity writes
@@
The relative importance between scattering and absorption is characterized by dust albedo . In the regime, it reduces to the case of the usual MBB emission in Eq. (11). For , the effect of dust scattering starts to become noticeable in the dust spectra at optical thick regime (Birnstiel et al. 2018). For most of the probable dust compositions (except for graphitic carbonaceous dust), this occurs when . First, dust scattering leads to attenuation of the observed intensity. Second, it changes the color (i.e., spectral indices), which depends on how the relative importance between and varies with frequency. Spectral indices are lowered (as compared to the dust with zero scattering opacity) in the frequency range where albedo is increasing with frequency, which was referred to as anomalous reddening (Liu 2019); conversely, spectral indices become larger in the frequency range where albedo is decreasing with frequency.
@@
To demonstrate how the hypothesis explain the observations, we generated isothermal track based on the dust grain model, of which the procedure is explained as follows. We first produced a grid of dust spectra based on various dust temperature, dust column density, and the maximum grain size . The lowest and highest temperatures are 6 K and 48 K, which correspond to the lower limit of typical interstellar medium (ISM) temperature and Rayleigh-Jeans limit, respectively. We then adopted the dust opacity tables provided in Birnstiel et al. (2018) which assume that dust grains are spherical and compact and are composed of water ice, refractory organics, troilite, and astrophysical silicates. For the grain size, we assumed a power-law grain size distribution with between the minimum grain size ( 0.1 m) and various . The size-averaged opacity is not sensitive to the assumption of . Last, surface densities start from to 30 g cm-2 to generate isothermal tracks from optically thin through thick cases, with a given and temperature. The Mie theory and the Henyey-Greenstein scattering approximation were considered.
@@
For micron-sized ( mm) dust grains, absorption dominates over scattering at millimeter bands. As is shown by the gray region in Fig. 10, the spectral index is close to the case of Eq. (16), so the highest possible spectral index is 3.8. As is discussed in Sect. 6.1, with mm, a large amount of sources with low spectral indices cannot be explained.
For mm-sized ( mm) dust grains, in the optically thin regime, the left and right boundaries of spectral indices shift toward higher values as compared to the case of mm. This is because albedo decreases with frequency at 100–200 GHz when mm (see Figures 2 and 3 of Liu 2019), which leads to higher spectral indices in the regime of . As is shown by the shaded pink region, the 1 mm sized grain scenario can cover the “high bump” that includes detections with . The high- bump feature actually serves as an observational hint for the existence of intermediate-sized grains in astrophysical environments. In the optically thick regime, the intensity attenuation is noticeable as compared with the case of mm, which better explains the data points with low and low .
@@
The origin of mm/cm-sized grains is uncertain. A typical envelope density of cm-3 is insufficient to form grains of this size. Only at densities of cm-3 would the formation of such large grains be feasible (Hirashita & Omukai 2009; Navarro-Almaida et al. 2024). According to coagulation models, mm-sized grains observed in protostellar envelopes cannot form locally within the envelopes themselves (Wong et al. 2016; Silsbee et al. 2022). In contrast to the inefficient dust growth in the envelope, dust can quickly grow to the order of mm in disks, with a typical timescale of years, thanks to higher densities in disks even in class 0/I YSOs. Numerical calculations have shown that large grains decoupled from gas can be ejected from disks through outflows due to the centrifugal force, enriching dust in envelopes (Wong et al. 2016; Tsukamoto et al. 2021). The timescale is much shorter than the lifetime of YSOs, allowing large grains to survive in envelopes.
@@ 6.2.3 Free-free emission
Massive protostars can ionize the surrounding gas, which produces bremsstrahlung (or free-free) emission by electronic braking. Free-free emission spectrum shows a weak frequency dependence in the optically thin regime at short wavelengths (Kurtz 2005), which can flatten the millimeter spectral index by MMB emission.
To explore the free-free emission contamination, we use the archived JVLA 1.3 cm observations with representative frequency of 23 GHz (Lu et al. 2019b) with rms noise of 0.05–0.15 mJy beam-1 as shown by yellow contours in Fig. 13. There exists extended filamentary 1.3 cm emission, which is suggested to have non-thermal origins (Clark & Caswell 1976; Ho et al. 1985; Lu et al. 2003; Meng et al. 2019; Bally et al. 2024). For example, the one at the northeast corner of the 20 km s-1 cloud has a spectral index of 0.4 (Ho et al. 1985; Padovani et al. 2019), which cannot be ascribed to free-free emission. In our analyses, two elongated structures (the aforementioned one in the 20 km s-1 cloud and one in cloud e) are excluded, because they have non-thermal origins.
At 23 GHz, the thermal free-free emission is optically thin even for the most compact HCHii regions (; turn-over frequency GHz) reported before (e.g., yang2019; yang2023). A spectral index of is assumed for the free-free emission in order to extrapolate from 1.3 cm to millimeter wavelength in the optically thin regime. We discuss the free-free contamination of all the associated sources in Appendix G.
However, these cases can only explain the low spectral indices of a limited number of sources that are associated with detectable radio emission. Below the 0.1 mJy sensitivity at 1.3 cm, there could still be some undetected UCHii regions. Below we quantitatively estimate how much the centimeter continuum contamination would be in order to significantly reduce the spectral index. Considering an optically thin source with a spectral index of 3.8, let the flux at 1.3 mm from pure dust emission be , the expected flux at 3 mm would then be . Taking the free-free emission flux at both 1.3 and 3 mm into account 777Free-free emission has % flux difference at 1.3 and 3 mm due to the weak frequency dependence ()., the differentiation of the spectral index can be rewritten from Eq. (1) as,
@@
Given that more than 70% of the sources exhibit low spectral indices, it is difficult to attribute them entirely to contamination by free-free emission. First, such a large number of massive protostars within a single cluster would conflict with expectations from the IMF, which predicts a decreasing population at the high-mass end. Second, low-mass protostars are unlikely able to contribute to the needed free-free emission either. In nearby molecular clouds (e.g., Corona Australis, Taurus, Ophiuchus, and Perseus; Liu et al. 2014; Dzib et al. 2015; Coutens et al. 2019; Tychoniec et al. 2018), free-free emission from low-mass class 0/I objects typically exhibits flux densities of only a few milli-Jansky or less at centimeter wavelengths. When scaled to the Galactic center, these emissions would become fainter by approximately three orders of magnitude, therefore contributing less than what could result in significant contamination at millimeter bands. To firmly exclude the possibility, the sensitivity needs to be at least (sub-)Jy. Indeed, a recent deep JVLA survey of the high-mass star-forming region G14.2250.50 ( 2 kpc) indicates that free-free emission from low-mass class 0/I YSOs in distant star-forming regions is challenging to detect with current facilities (Díaz-Márquez et al. 2024).